Trying to solve this equation without success:
(e^x)-(e^(-x))=12

Answer by stanbon(75887)

Trying to solve this equation without success:
(e^x)-(e^(-x))=12
------
e^x - (1/e^x) = 12
----
e^(2x) - 1 = 12e^x
----
Quadratic Form::
(e^x)^2 - 12(e^x) - 1 = 0
------
Solve for e^x ::
e^x = 12.08276
---
x = ln(12.08276)
x = 2.49178..
Cheers,
Stan H.
-----------